What are the zeros of f(x) = 2x2 - 2x -220?
Answer #1

Given,

f(x) = 2x2-21-220

To find the zeros of f(x)

Equate f(x) to 0

2x^{2}-2x-220=0

Dividing with 2 on both sides

x^{2}-x-110=0

x^{2}-11x+10x-110=0

x(x-11)+10(x-11)=0

(x+10)(x-11)=0

(x+10)=0,(x-11)=0

x=-10,x=11

x=-10,11

Therefore,

The zeros of f(x) are x=-10,11